OpenStudy (anonymous):

The average middle-distance runner at a local high school runs the mile in 4.5 minutes, with a standard deviation of 0.3 minute. Find the percentage of a runners that will run the mile in less than 4 minutes? A. 45.22% B. 15.87% C. 7.15% D. 4.75%

OpenStudy (anonymous):

P(x<4) =P((x-4.5)/0.3<(4-4.5)/0.3) =P(z<-1.6666)=1-\[\phi\](1.6666)=0.1587

OpenStudy (anonymous):

thanks!